OpenStudy (anonymous):

Function Notations Given f(x) = 2x-7 g(x) = x^2 + x Find: g(f(x))- I think I have the wrong answer, that is probably why I am confused out of my mind :S. Please help me with full steps. :)

OpenStudy (anonymous):

g(f(x))=(f(x))^2+f(x)= (2x-7)^2+2x-7=4x^2-26x+42

OpenStudy (anonymous):

\[f(x) = 2x - 7\] \[g(x) = x^2 + x\] Substituting \[g(f(x)) = (2x - 7)^2 + (2x - 7)\] Expanding \[g(f(x)) = 4x^2 - 28x + 49 + 2x - 7\] Simplifying \[g(f(x)) = 4x^2 - 26x + 42\] :)